When a hydrogen atom emits a photon of energy 12.1ev, it's orbital angular momentum changes by_ _ _

Grade:12th pass

1 Answers

Eshan
5 years ago
Dear student,

12.1eV energy is emitted when electron makes a transition from n=3 to n=1.

Angular momentum of electrons in Bohr orbit is given as\dfrac{nh}{2\pi}.

Hence change in angular momentum=\dfrac{(3-1)h}{2\pi}=\dfrac{h}{\pi}